OpenStudy (anonymous):

Kofi estimated that he would run a total of 10 km in 1 h. In 1 h, he actually ran 46 times around a track 0.25 km long, going a total distance of 11.5 km. What was the percent error in Kofi’s estimate? Round to the nearest tenth of a percent, if necessary. A. 1.5% B. 5% C. 13% D. 23%

OpenStudy (anonymous):

@Le_Asien

OpenStudy (anonymous):

well, this was slightly complicated but I believe I got it

OpenStudy (anonymous):

100/10= 10 10*11.5=115 100/115=0.87 1-0.87=.13 0.13= 13 % Got It?

OpenStudy (anonymous):

happy with the answer?